OpenStudy (anonymous):

can someone help me with expansion and give me like practise problems plssss????

OpenStudy (anonymous):

You mean multiplication of algebraic expressions?

OpenStudy (anonymous):

no exspansion of algebraic epression

OpenStudy (callisto):

Try these perhaps :P 1.\[ 3x^2 y^4 (2x-3y+4z)\] 2. \[3\sqrt{x} (x^4 -x^2y^5)\]

OpenStudy (anonymous):

@Callisto that's too difficult for him, give him something easier... \[\Large x(3x^2+4)=\] \[\Large x^5(3x^2-2x)=\] then try callisto's one ! ;)

OpenStudy (callisto):

or these... 1. 2(x+y) 2. 12 ( x^2 + xy^2 - x^3)

OpenStudy (callisto):

@Kreshnik my new ones are easier now :)

OpenStudy (anonymous):

@Callisto I agree ! ;)

OpenStudy (callisto):

First, explain a little for you.. 2(x+y) = 2(x) + 2(y) = 2x+2y Can you try the 2nd question of my 2nd post?

OpenStudy (anonymous):

\[\Large x^2(2x+4)=2x\cdot x^2+4\cdot x^2\] \[\Large x^4(2x^2-7x)=x^4\cdot 2x^2-x^4\cdot 7x=2x^{4+2}-7x^{4+1}=2x^6-7x^5\] Exampless...

OpenStudy (anonymous):

\[\Large 2x\cdot x^2+4x^2=?\] and you also are supposed to know these rules ! : \[\LARGE x^a\cdot x^b=x^{a+b}\] \[\LARGE \frac{x^a}{x^b}=x^{a-b}\] \[\LARGE x^{-a}=\frac{1}{x^a} \quad \quad and \quad \quad \frac{1}{x^a}=x^{-a}\] \[\LARGE \sqrt[b]{x^a}=\left(x\right)^{\frac ab}\]